<e> Solve using the method of separation of variables the pde partial du/dt+du/dx+2e^tu=0


1
Expert's answer
2022-02-18T13:07:49-0500

By the method of separation of variables we suppose that "u" is of the form "u(t,x)=\phi(t)\psi(x)" which gives us :

"\phi'(t) \psi(x)+\phi(t)\psi'(x)+2e^t \phi(t) \psi(x)=0"

dividing both parts by "u(t,x)=\phi(t)\psi(x)" gives us

"\phi'(t)/\phi(t) + 2e^t+\psi'(x)/\psi(x)=0"

"\phi'(t)/\phi(t)+2e^t=-\psi'(x)/\psi(x)"

As the left part is a function of "t" and the right is the function of "x" the equality gives us

"\begin{cases} \phi'/\phi +2e^t=\lambda \\ -\psi'/\psi = \lambda \end{cases}" with "\lambda\in \mathbb{R}" a constant.

The second equation easily gives us "\psi_\lambda (x)= Ae^{-\lambda x}", "A" a constant

The first one gives us

"\ln \phi(t) - \ln\phi(0)=\lambda t -2e^t+2"

"\phi_\lambda(t)=Be^{\lambda t - 2e^t}" , "B" a constant

Combining two solutions gives us

"u_\lambda (t,x)= Ce^{\lambda(t-x)} e^{-2e^t}"